an iron ring having area 400mm² is mean diameter 25cm having number of turns 500.Assume relative permeability 250 resistance 474 ohm and voltage 240v . Find the flux.
1
Expert's answer
2019-11-11T15:57:29-0500

For toroid


"B=\mu\mu_0\frac{NI}{\pi d}=\mu\mu_0\frac{NU}{\pi dR}"


So, we have


"\Phi_c=BS\cos\alpha=\mu\mu_0\frac{NU}{\pi dR}S"


Finally


"\Psi=NB=\mu\mu_0\frac{N^2 U}{\pi dR}S=250\cdot4\cdot3.14\cdot10^{-7} \cdot\frac{500^2\cdot240}{3.14\cdot0.25\cdot474}\cdot400\cdot10^{-6 }=0.02Wb"


Best